OpenStudy (anonymous):

Can someone help me make an example and complete it of Heron's formula for finding the area of a triangle?

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

pick three numbers make them small and easy to work with

OpenStudy (anonymous):

i can pick if you like... i would pick \(3,4,5\)

OpenStudy (anonymous):

Sounds good to me!

OpenStudy (anonymous):

kind of cheating, because this is a a right triangle, so the answer should be \(\frac{1}{2}\times 3\times 4=6\) but we can do it anyway

OpenStudy (anonymous):

first you need \[s=\frac{a+b+c}{2}=\frac{3+4+5}{2}=6\]

OpenStudy (anonymous):

then you compute \[\sqrt{s(s-a)(s-b)(s-3)}\] \[=\sqrt{6(6-3)(6-4)(6-5)}\] \[=\sqrt{6\times 3\times 2\times 1}\] \[=\sqrt{36}\] \[=6\] as predicted

OpenStudy (anonymous):

Okay..

OpenStudy (anonymous):

Thank you so much for your help.